OpenStudy (anonymous):

The sum of three consecutive integers is 72. Find the integers. (2 points) 22, 23, 24 25, 26, 27 23, 24, 25 24, 25, 26

OpenStudy (anonymous):

23,24,25 just add them all.

OpenStudy (mertsj):

Let the integers be: x= first integer x+1 = second integer x+2=third integer Then add them up and you should get 72 x+(x+1)+(x+2)=72 Can you solve that equation?

OpenStudy (anonymous):

As the integers are consecutive, you can take the three consecutive integers as x,(x+1),(x+2)

OpenStudy (anonymous):

yea i can solve that do u want me too?

OpenStudy (mertsj):

Yes. Please

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

The quickest way to find this is just check the options, but the analytically it can be solved as shown by mertj.

OpenStudy (anonymous):

okay i think it would be like this 1x +1x+1+1x+2=72 ?

OpenStudy (mertsj):

Yes. Add the like terms.

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

would it be 3x+3 ?

OpenStudy (mertsj):

It would be. Very good. So now we have: 3x+3=72

OpenStudy (mertsj):

Can you solve that?

OpenStudy (anonymous):

yes...hold on...

OpenStudy (anonymous):

it would be x=23

OpenStudy (mertsj):

Excellent. Now we have to be sure to answer the question the problem asks. So reread the question and answer it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it would be 23,24,25 ?

OpenStudy (mertsj):

Good for you. Very smart girl!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Let the 3 integers be: x, x+1, x+2 Given that: x+x+1+x+2=72 3x+3=72 3(x+1)=72 x+1=72/3 x+1=24 x=24-1 x=23 So, x+1=24, and x+2=25 So the choices are 23,24, and 25. Which is choice C. or 3.

OpenStudy (mertsj):

You're welcome. Study hard. You will be successful.

OpenStudy (anonymous):

okay thats 4 ur time:).. happy valentines day

OpenStudy (mertsj):

You too. I hope your honey gets you a nice box of candy and flowers.